40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1296     Entry time: Thu Feb 12 11:21:54 2009
Author: Yoichi 
Type: Update 
Category: LSC 
Subject: Locking effort resumed 
Last night, I restarted the locking work.
Quite some time was wasted by the disconnected REFL199 by Alberto for the cavity length measurement.
From now on, please put the interferometer back to the original state every day.
If possible, please refrain from changing the IFO settings (cabling, optics, etc).
It is also very important to always restore the full IFO alignment after you are done with your work.

While I was working on the optimization of the DD hand-off, the DRMI alignment got into a strange state.
Even when I did the whole dither alignment procedure from the beginning (from x-arm), the AS166Q did not go above 1000.
PRMI looks ok (SPOB goes above 1100). I could lock the DRMI but the lock position hops to other modes easily.
Manual tweaks of SRM did not help.
After running the whole alignment procedure several times in vain, I was too tired and went home.
I noticed that the single arm lock shows power drops again. There are some offsets in the arm lock loops.
This may have prevented the Michelson alignment from being optimal. I will check this today.
ELOG V3.1.3-